HanAkoSolo (jamierox4ev3r):

Reviewing some math basics before I start college. I'm unsure of a few things, so please check my work as I go. Thank you. First question: 2c. Simplify the following expression. Write your answer without negative exponents.

HanAkoSolo (jamierox4ev3r):

Give me some time to type out the expression, thanks.

HanAkoSolo (jamierox4ev3r):

\[\left( \frac{ 3x ^{\frac{ 3 }{ 2 }}y ^{3} }{ x ^{2} y ^{-\frac{ 1 }{ 2 }}} \right)^{-2}\]

HanAkoSolo (jamierox4ev3r):

oh my goodness that was a heck of an expression to get out. So far, what I've done is this: \[\left( \frac{ x ^{2}y ^{-\frac{ 1 }{ 2 }} }{ 3x ^{\frac{ 3 }{ 2 }} y ^{3}} \right)^{2}\] basically, I got ride of the negative 2 exponent by flipping the fraction around. If this was an incorrect thing to do, someone let me know now. My recollection of algebra rules is a bit foggy

HanAkoSolo (jamierox4ev3r):

any idea? @Nnesha

Nnesha (nnesha):

wait lol :P

HanAkoSolo (jamierox4ev3r):

alright :P thanks

Nnesha (nnesha):

alright back

Nnesha (nnesha):

sorry abt that i need to buy a new laptop >.< i'm just gonna use draw tool  but i'm sure we will get the same answer :=) yes